OpenStudy (anonymous):

Solve 2x + 5y = 10 for y.

OpenStudy (anonymous):

y=(10-2x)/5

OpenStudy (anonymous):

\[5y = 10-2x => y=\frac{10-2x}{5}\]

OpenStudy (anonymous):

\[2x+5y=10\] \[5y=10-2x\] \[y=\frac{10-2x}5\]